# The Dial-a-Ride Problem with Limited Pickups per Trip

### v2

@@ 3.3 Fragment-based formulations of DARP-LPT
The framework presented by Rist & Forbes (2021) defines a fragment-based formulation for DARP, which necessitates the inclusion of callbacks to handle rarely violated constraints related to subtour elimination, time windows, and maximum ride time. In the case of DARP, even when multiple vehicles are used, this fragment-based format remains feasible. In this section, our fragment-based formulations build upon the work of Rist & Forbes (2021). However, since DARP-LPT involves frequently violated constraints, in particular the limitations on the number of pickups per trip (per vehicle), we introduce different decision variables, shifting from a fragment-based format to a fragment-vehicle-based format. Consequently, we propose two distinct fragment-based formulations tailored for multi-vehicle settings: a fragment flow formulation and a fragment assignment formulation. The former assigns a vehicle to each fragment and node arc, while the latter exclusively links a vehicle with each fragment. In particular, the latter formulation draws inspiration from the client-depot assignment formulation in multi-depot routing problems (Bektaş et al., 2020). In our context, the “client” corresponds to the fragment, while the “depot” corresponds to the vehicle.
@@ 3.3.1 Preliminaries and notation
In this subsection, we largely follow the definitions of Rist & Forbes (2021).
@@
Rist & Forbes (2021) introduce the concept of restricted fragments for DARP. These fragments can be combined to enumerate all DARP routes, and a network is constructed to include all these restricted fragments. Prior to introducing restricted fragments, we describe a full fragment in more detail.
@@
A restricted fragment (RF) is a part of a full fragment, starting at a pickup node whose predecessor is a delivery node or depot and ending at a delivery node whose successor is a pickup node or depot. An RF contains at least one pickup-to-delivery movement (not necessarily corresponding to a customer’s pickup and delivery location pair) and cannot be divided into smaller RFs (meaning each RF is as small as possible), e.g., $(p1,p2,d1)$, $(p1,d2,d1)$, $(p1,p2,p3,d4,d5)$, $(p5,d8)$, $(p5,p7,d1,d2,d3)$. The following two examples are not RFs: $(p1,p2,p3,p4)$ and $(p2,d1,p3,d3,d2)$, where the former lacks a pickup-to-delivery movement and the latter can be split after $d1$. A special type of RF is a FF with the format “$ppppdddd$”, where all pickups precede deliveries, and which cannot be divided into smaller RFs, for instance $(p1,p2,d1,d2)$ or $(p1,p3,p2,d2,d1,d3)$. Below are two further examples, where the left-hand side of the equality is the route path of an FF, and the right-hand side consists of RFs. This decomposition is entirely unique.
@@
A node is a pair consisting of a location and a set of loads (with customer identification and not just the number of customers) on board, denoted as $(loc(i),loadset(i))$ (or just $loc(i)$ if $loadset(i)=\emptyset$). The $loadset(i)$ includes the unserved customers along a path after departing from location $loc(i)$ but excludes those from $loc(i)$. For instance, the nodes from the route path $(p1,p2,d1,p3,d2,p4,d3,d4)$ are expressed as $p1$, $(p2,\{l_{1}\})$, $(d1,\{l_{2}\})$, $(p3,\{l_{2}\})$, $(d2,\{l_{3}\})$, $(p4,\{l_{3}\})$, $(d3,\{l_{4}\})$, and $d4$. Notably, nodes with the same location can be distinct, as illustrated by examples such as $(p2,\{l_{1}\})$, $(p2,\{l_{3}\})$, $(p2,\{l_{2},l_{3}\})$, and $p2$. We define a node with a pickup location as a “pickup node” and a node with a delivery location as a “delivery node”. Specifically, we denote the sets of pickup and delivery nodes as $P_{N}$ and $D_{N}$, respectively, distinguishing them from the sets of pickup and delivery locations denoted as $P$ and $D$.
@@ 3.3.2 Fragment-based network
Building upon the introductory concepts, we now introduce the fragment-based network. To facilitate comprehension, we present several figures (Fig. 1, Fig. 2, and Fig. 3). In these figures, ellipses represent nodes, solid lines (with ellipses on both sides) indicate fragments (with their route paths written above or to the right), and dashed lines with arrows are node arcs (further explained below). $O^{+}$ and $O^{-}$ denote the origin depot and destination depot, respectively.
@@
Decomposing feasible DARP routes into fragments yields a significantly smaller number of fragments compared to DARP routes. Conversely, if we combine these fragments, we can obtain all feasible DARP routes, albeit with the possibility of some redundant routes. Fig. 2 illustrates how feasible DARP routes can be gathered in a transitional network (as an intermediate step towards the fragment-based network). Fig. 1 presents several DARP routes and their corresponding fragments. The network in Fig. 2 will use these fragments and allow to enumerate all the DARP routes shown in Fig. 1. Identical fragments across different routes in Fig. 1 can be consolidated into a single fragment while maintaining the connections to other fragments. Subsequently, the fragment-based network of Fig. 3 is derived from the network of Fig. 2. When different fragments share the same start and/or end nodes, then those nodes can also be merged. For instance, two fragments with the same start and end nodes but different route paths $(p1,d1)$ and $(p1,p4,d4,d1)$ become two route paths originating from node $p1$ and terminating at node $d1$. Such consolidations help reduce connections (node arcs). Node arcs with the same start node and end node can also be merged; the three connections from $(p1,d1)$ to $(p2,d2)$, from $(p1,d1)$ to $(p2,p3,d2,d3)$, and from $(p1,p4,d4,d1)$ to $(p2,d2)$ in Fig. 2, for instance, can be merged into one node arc $(d1,p2)$ in Fig. 3.
Figure 2: A transitional network for DARP by merging all identical fragments from Fig. 1
[…]
Figure 3: Fragment-based network for DARP by merging all identical nodes from Fig. 2
The practical process of network generation proceeds as follows. The fragment-based network comprises nodes, fragments, and node arcs. The initial step involves generating fragments (outlined in Section 4). Following this, the start and end nodes of all fragments are identified and added to the network. Subsequently, all fragments are integrated into the network, connecting from their start nodes to end nodes. Different fragments can share the same start and/or end nodes. In the final phase, for the node arcs, all delivery nodes and the origin depot connect with all pickup nodes, and all delivery nodes connect with the destination depot. If none of the fragments ending at $h^{-}$ can connect to any fragment starting at $h^{+}$ due to time window or any other constraints, the node arc $(h^{-},h^{+})$ can be eliminated. When the network exclusively includes FFs, the number of node arcs (excluding those involving depots) is at most $|n|\times|n|$. For the network with RFs, the potential number of node arcs could be higher due to the possibility of different nodes with the same location having different loads. Moreover, a feasible route typically starts with a node arc from the depot, followed by a fragment leading to a delivery node. Subsequently, there must be a node arc from the current delivery node to another pickup node or the destination depot. After that, if the current node is not the depot yet, it must be succeeded by another fragment and then a node arc again, continuing in this pattern. Therefore, any feasible route in Fig. 3 must adhere to the alternating pattern of dashed lines and solid lines.
@@ 3.3.3 Fragment flow formulation
The first fragment-based formulation is the fragment flow formulation (FFF), which expands on the framework of Rist & Forbes (2021). It transitions from a formulation using fragment-based and node arc-based decision variables to one employing fragment-vehicle-based and node arc-vehicle-based variables. Please note that the current formulation below is not complete. Lazy constraints, specifically designed for the elimination of infeasible paths, are applied to address subtour elimination, time windows, and maximum ride time. The details are described in Section 3.3.6.
@@
3.3.4 Fragment assignment formulation
Table 3: Decision variables of FAF
@@ 3.3.4 Fragment assignment formulation
$X_{f}$ = 1, if fragment $f\in F$ is traversed; = 0, otherwise
$Y_{a}$ = 1, if node arc $a\in A_{N}$ is traversed; = 0, otherwise
An alternative formulation involves the assignment of fragments to vehicles and is called fragment assignment formulation (FAF). FAF extends the framework of Rist & Forbes (2021) by transitioning from a formulation using fragment-based decision variables to one that incorporates fragment-vehicle-based variables. Unlike FFF, however, which introduces node arc-vehicle-based decision variables, FAF keeps the node arc-based variables by means of a separate set of constraints (which hinges upon constraint set (31)). This model is inspired by the client-depot assignment formulation used in multi-depot routing problems (Bektaş et al., 2020). In the same way as for FFF, the current formulation below is not complete. Lazy constraints to address subtour elimination, time windows, and maximum ride time are discussed in Section 3.3.6.
$\min$ $\sum_{f\in F}c_{f}X_{f}+\sum_{a\in A_{N}}c_{a}Y_{a}$ (24)
$s.t.$ $\sum_{f\in F_{h}^{-}}X_{f}$ $=\sum_{a\in A_{h}^{+}}Y_{a}$ $h\in P_{N}$ (25)
$\sum_{f\in F_{h}^{+}}X_{f}$ $=\sum_{a\in A_{h}^{-}}Y_{a}$ $h\in D_{N}$ (26)
$\sum_{f\in F_{i}}X_{f}$ $=1$ $i\in P$ (27)
$\sum_{a\in A_{O^{+}}^{+}}Y_{a}$ $\leq|V|$ (28)
$x_{fv}$ $\leq X_{f}$ $f\in F,v\in V$ (29)
$\sum_{v\in V}x_{fv}$ $=X_{f}$ $f\in F$ (30)
$\sum_{f\in F^{-}_{i}}x_{fv}+Y_{a}$ $\leq\sum_{f\in F^{+}_{j}}x_{fv}+1$ $a=(i,j)\in A_{N}:i\neq O^{+},j\neq O^{-},v\in V$ (31)
$\sum_{f\in F}L_{f}x_{fv}$ $\leq L$ $v\in V$ (32)
$X_{f},Y_{a}$$\in\{0,1\}$ $f{\in}F,a{\in}A_{N}$ (33)
$x_{fv}$$\in\{0,1\}$ $f{\in}F,v{\in}V$ (34)
The objective function (24) minimizes the weighted cost of used fragments and node arcs. Constraints (25) and (26) describe the network flow for fragments and node arcs. Constraints (27) guarantee that each pickup location is visited exactly once. Constraints (28) state that at most $|V|$ vehicles are used. Constraints (29)-(31) assign the fragments to vehicles. Specifically, constraints (29) impose that if a fragment is traversed by a vehicle, the fragment is used; constraints (30) state that if a fragment is chosen, it must be assigned to a vehicle; and constraints (31) express that if a node arc is used, the fragments preceding the arc and the fragments succeeding the arc must be traversed by the same vehicle. This explains why it does not matter which vehicle traverses the node arc. Constraints (32) set the maximal number of pickups. Constraints (33) and constraints (34) specify the domains of $X$, $Y$, and $x$.
Although the variables $X$ and $x$ in this formulation can be expressed solely in terms of $x$, keeping both $X$ and $x$ yields better computational results.
3.3.5 Comparison of FFF and FAF
In this subsection, we will establish that FFF is a stronger formulation than FAF. We will show this in two steps. Firstly, any feasible solution to the linear relaxation of FFF also qualifies as a feasible solution to the linear relaxation of FAF. Secondly, we will provide an example solution from the polytope of FAF that does not appear in the polytope of FFF.
Regarding the first step, it is easy to observe that constraints (18)-(20) and constraints (21) in FFF can imply constraints (25)-(27) and constraints (32) in FAF. Furthermore, constraints (23) in FFF can indicate constraints (28) in FAF. Therefore, our only task is to demonstrate how a solution ($x^{\prime}$, $y^{\prime}$) that is feasible under FFF constraints (18) and (19) can meet the requirements of FAF constraints (31). Let ($x^{\prime}$, $y^{\prime}$) represent a solution from the linear relaxation of FFF. We introduce two auxiliary variables: $X^{\prime}$ and $Y^{\prime}$. Here, $X^{\prime}_{f}=\sum_{v\in V}x_{fv}$ for all $f\in F$ and $v\in V$; and $Y^{\prime}_{f}=\sum_{v\in V}y^{\prime}_{av}$ for all $a\in A_{N}$ and $v\in V$.
Proposition 1.
Constraints (18) and constraints (19) in FFF imply constraints (31) in FAF.
Proof.
Consider an arc $a=(i,j)$. Our objective is to demonstrate that any solution $(x^{\prime},y^{\prime},X^{\prime},Y^{\prime})$ satisfying the FFF constraints will also be feasible for the FAF network flow constraints (31).
According to the FFF network flow constraints (18) and (19), we have:
$\sum_{f\in F_{i}^{-}}x^{\prime}_{fv}=\sum_{a\in A_{i}^{+}}y^{\prime}_{av},\quad\text{and}\quad\sum_{f\in F_{j}^{+}}x^{\prime}_{fv}=\sum_{a\in A_{j}^{-}}y^{\prime}_{av}$
Thus, we can transform the FAF constraints (31) as follows:
$\sum_{f\in F^{-}_{i}}x^{\prime}_{fv}+Y^{\prime}_{a}\leq\sum_{f\in F^{+}_{j}}x^{\prime}_{fv}+1\iff\sum_{a^{\prime}\in A_{i}^{+}}y^{\prime}_{a^{\prime}v}+Y^{\prime}_{a}\leq\sum_{a^{\prime}\in A_{j}^{-}}y^{\prime}_{a^{\prime}v}+1\iff\sum_{a^{\prime}\in A_{i}^{+}\setminus a}y^{\prime}_{a^{\prime}v}+Y^{\prime}_{a}\leq\sum_{a^{\prime}\in A_{j}^{-}\setminus a}y^{\prime}_{a^{\prime}v}+1$
We now examine the left side of the last inequality, $\sum_{a^{\prime}\in A_{i}^{+}\setminus a}y^{\prime}_{a^{\prime}v}+Y^{\prime}_{a}$, using FFF constraints (18). Then, we obtain:
$\sum_{a^{\prime}\in A_{i}^{+}\setminus a}y^{\prime}_{a^{\prime}v}+Y^{\prime}_{a}\leq\sum_{a\in A_{i}^{+}}Y_{a}=\sum_{f\in F^{-}_{i}}X_{f}\leq 1\leq\sum_{a^{\prime}\in A_{j}^{-}\setminus a}y^{\prime}_{a^{\prime}v}+1$
Thus, we see that constraints (31) are automatically fulfilled if constraints (18) and (19) hold. ∎
The second step is to show that the polyhedra of the two formulations do not coincide. To illustrate this point, we provide an example solution of FAF that is cut away by FFF. For example: if $V=\{v1,v2\}$, $F^{-}_{i}=\{f\}$, $F^{+}_{j}=\{g\}$, and $a=(i,j)$, constraints (31) of FAF become: $x_{fv1}+x_{fv2}+Y_{a}\leq x_{gv1}+x_{gv2}+1$. A potential solution under the given constraints is $x_{fv1}=0.5$, $x_{fv2}=0.5$, $Y_{a}=1$, $x_{gv1}=1$, $x_{gv2}=0$. However, this solution does not satisfy constraints (18) and (19) of FFF, where $x_{fv2}=y_{av2}$ and $y_{av2}=x_{gv2}$.
Based on the foregoing, we conclude that FFF dominates FAF. Despite this dominance result, we should still verify empirically how the two formulations behave in terms of runtimes. FAF, for instance, has less decision variables since FAF utilizes arc-based variables $Y$, whereas FFF uses arc-vehicle-based variables $y$.

### v3

@@ 3.3 Fragment-based formulations of DARP-LPT
The framework presented by Rist & Forbes [18] defines a fragment-based formulation for DARP, which necessitates the inclusion of callbacks to handle rarely violated constraints related to subtour elimination, time windows, and maximum ride time. In the case of DARP, even when multiple vehicles are used, this fragment-based format remains feasible.
This section builds upon the fragment-based formulations introduced by Rist & Forbes [18]. We propose two extended formulations to suitably manage the frequent violation of specific constraints in DARP-LPT, particularly the limited number of pickups per trip. Formulation FFF augments the decision variables by incorporating vehicle indices, while PSFF reformulates the network using a pickup-space structure and omits vehicle details. In the case of MDARP-LPT, FFF retains its structure, whereas PSFF must reintroduce vehicle indices to account for the unique origin and destination associated with each vehicle.
@@ 3.3.1 Preliminaries and notation
In this subsection, we largely follow the definitions of Rist & Forbes [18].
@@
Rist & Forbes [18] introduce the concept of restricted fragments for DARP. These fragments can be combined to enumerate all DARP routes, and a network is constructed to include all these restricted fragments. Prior to introducing restricted fragments, we describe a full fragment in more detail.
@@
A restricted fragment (RF) is a part of a full fragment, starting at a pickup node whose predecessor is a delivery node or depot and ending at a delivery node whose successor is a pickup node or depot. An RF contains exactly one pickup-to-delivery movement (not necessarily corresponding to a customer’s pickup and delivery location pair), e.g., $(p1,p2,d1)$, $(p1,d2,d1)$, $(p1,p2,p3,d4,d5)$, $(p5,d8)$, $(p5,p7,d1,d2,d3)$. The following two examples are not RFs: $(p1,p2,p3,p4)$ and $(p2,d1,p3,d3,d2)$, where the former lacks a pickup-to-delivery movement and the latter can be split after $d1$. A special type of RF is a FF with the format “$ppppdddd$”, where all pickups precede deliveries, and which cannot be divided into smaller RFs, for instance $(p1,p2,d1,d2)$ or $(p1,p3,p2,d2,d1,d3)$. Below are two further examples, where the left-hand side of the equality is the route path of an FF, and the right-hand side consists of RFs. This decomposition is entirely unique.
@@
A node is a pair consisting of a location and a set of loads (with customer identification and not just the number of customers) on board, denoted as $(loc(i),loadset(i))$ (or just $loc(i)$ if $loadset(i)=\emptyset$). The $loadset(i)$ includes the unserved customers along a path after departing from location $loc(i)$ but excludes those from $loc(i)$. For instance, the nodes from the route path $(p1,p2,d1,p3,d2,p4,d3,d4)$ are expressed as $p1$, $(p2,\{l_{1}\})$, $(d1,\{l_{2}\})$, $(p3,\{l_{2}\})$, $(d2,\{l_{3}\})$, $(p4,\{l_{3}\})$, $(d3,\{l_{4}\})$, and $d4$. Notably, nodes with the same location can be distinct, as illustrated by examples such as $(p2,\{l_{1}\})$, $(p2,\{l_{3}\})$, $(p2,\{l_{1},l_{3}\})$, and $p2$. We define a node with a pickup location as a “pickup node” and a node with a delivery location as a “delivery node”. Specifically, we denote the sets of pickup and delivery nodes as $P_{N}$ and $D_{N}$, respectively, distinguishing them from the sets of pickup and delivery locations denoted as $P$ and $D$.
@@ 3.3.2 Fragment-based network
Building upon the introductory concepts, we now introduce the fragment-based network. To facilitate comprehension, we provide Fig. 1 and Fig. 2. Fig. 1 illustrates several DARP routes alongside their corresponding fragments, whereas Fig. 2 depicts the corresponding fragment-based network. This network can be derived from Fig. 1 by first merging all identical fragments (with an intermediate transformation shown in Fig. A.1 in A), and then merging identical nodes while preserving the original connections. In these figures, ellipses represent nodes, solid lines (with ellipses on both sides) indicate fragments (with their route paths written above or to the right), and dashed lines with arrows are node arcs (further explained below). $O^{+}$ and $O^{-}$ denote the origin depot and destination depot, respectively.
@@
[…]
Figure 2: Fragment-based network for DARP by merging all identical fragments and nodes from Fig. 1
@@
The practical process of network generation proceeds as follows. The fragment-based network comprises nodes, fragments, and node arcs. The initial step involves generating fragments (outlined in Section 4). Following this, the start and end nodes of all fragments are identified and added to the network. Subsequently, all fragments are integrated into the network, connecting from their start nodes to end nodes. Different fragments can share the same start and/or end nodes. In the final phase, for the node arcs, all delivery nodes and the origin depot connect with all pickup nodes, and all delivery nodes connect with the destination depot. If none of the fragments ending at $h^{-}$ can connect to any fragment starting at $h^{+}$ due to time window or any other constraints, the node arc $(h^{-},h^{+})$ can be eliminated. When the network exclusively includes FFs, the number of node arcs (excluding those involving depots) is at most $|n|\times|n|$. For the network with RFs, the potential number of node arcs could be higher due to the possibility of different nodes with the same location having different loads. Moreover, a feasible route typically starts with a node arc from the depot, followed by a fragment leading to a delivery node. Subsequently, there must be a node arc from the current delivery node to another pickup node or the destination depot. After that, if the current node is not the depot yet, it must be succeeded by another fragment and then a node arc again, continuing in this pattern. Therefore, any feasible route in Fig. 2 must adhere to the alternating pattern of dashed lines and solid lines. For details on why Fig. 2 can be used to enumerate all paths, please refer to A.
@@ 3.3.3 Fragment flow formulation
The first fragment-based formulation is the fragment flow formulation (FFF), which expands on the framework of Rist & Forbes [18]. It transitions from a formulation using fragment-based and node arc-based decision variables to one employing fragment-vehicle-based and node arc-vehicle-based variables. Please note that the current formulation below is not complete. Lazy constraints, specifically designed for the elimination of infeasible paths, are applied to address subtour elimination, time windows, and maximum ride time. The details are described in Section 3.3.5.
@@
Furthermore, to enhance the formulation FFF, we introduce additional constraints based on the minimum number of used vehicles, as shown in constraints (24). While these constraints may be redundant, they significantly speed up the computation process. Additionally, for constraints (23), we set $w_{v}$ equal to 1 for $v\in V_{f}$.
$\sum_{f\in F}x_{vf}$ $\geq 1$ $v\in V_{f}$ (24)
3.3.4 Pickup-space fragment formulation
We now describe the Pickup-Space Fragment Formulation (PSFF), which extends the fragment-based formulation of Rist & Forbes [18] by introducing the number of pickups per trip as a new dimension, analogous to a time-space representation. Unlike FFF, which uses a separate constraint for pickup limits, PSFF integrates this constraint directly into the pickup-space network.
The pickup-space network is formed by replicating the fragment-based network across $L+1$ levels, where each level represents the remaining number of allowable pickups throughout a trip. The index sets $\mathcal{L_{P}}=\{1,\dots,L\}$ and $\mathcal{L_{D}}=\{0,\dots,L-1\}$ define the remaining pickup levels for pickup and delivery nodes, respectively. When a vehicle traverses a fragment with $l\in\mathcal{L_{P}}$ remaining allowable pickups before entry, it concludes at level $l-L_{f}\in\mathcal{L_{D}}$. The transitions for node arcs adhere to the same level structure, ensuring consistency in tracking the remaining allowable number of pickups.
The following formulation does not include all constraints. Lazy constraints are incorporated to address subtour elimination, time windows, and ride-time constraints, as detailed in Section 3.3.5.
Table 3: Decision variables of LSF
@@ 3.3.4 Pickup-space fragment formulation
$X_{fl}$ = 1, if a vehicle traverses fragment $f\in F$ with $l$ remaining allowable pickups
before entry; = 0, otherwise
$Y_{al}$ = 1, if a vehicle traverses node arc $a\in A_{N}$ while having $l$ remaining allowable
pickups available; = 0, otherwise
$\min\sum_{l\in\mathcal{L_{P}}}\sum_{f\in F}c_{f}X_{fl}+\sum_{l\in\mathcal{L_{D}}}\sum_{a\in A_{N}}c_{a}Y_{al}$ (25)
$s.t.\quad$ $\sum_{f\in F_{h}^{-}}X_{f,l+L_{f}}$ $=\sum_{a\in A_{h}^{+}}Y_{al}$ $h\in P_{N},\;l\in\mathcal{L_{P}}$ (26)
$\sum_{f\in F_{h}^{+}}X_{fl}$ $=\sum_{a\in A_{h}^{-}}Y_{al}$ $h\in D_{N},\;l\in\mathcal{L_{D}}$ (27)
$\sum_{l\in\mathcal{L_{P}}}\sum_{f\in F_{i}}X_{fl}$ $=1$ $i\in P_{N}$ (28)
$\sum_{a\in A_{o}^{+}}Y_{aL}$ $\leq|V|$ (29)
$X_{fl}$$\in\{0,1\}$ $f\in F\;l\in\mathcal{L_{P}}$ (30)
$Y_{al}$$\in\{0,1\}$ $a\in A_{N}\;l\in\mathcal{L_{D}}$ (31)
The objective 25 is to minimize the total weighted cost of fragments and arcs. Constraints (26) and (27) establish the network flow constraints on each pickup and delivery node at every remaining pickup level. These constraints help track the remaining allowable pickup number throughout the trip. Specifically, constraints (26) ensure that if a fragment $f$ initially has a remaining allowable pickup number of $l+L_{f}$ then it will end at a delivery node with level $l$, and then a node arc must depart from this delivery node at the same level. Constraints (27) specify that if a fragment departs from a pickup node at level $l$, then a node arc must end at this pickup node at level $l$ (and this node arc also starts at level $l$). Constraint (28) ensures that each pickup location is visited exactly once, thereby guaranteeing that all customer requests are fulfilled. Constraint (29) imposes an upper bound on the total number of vehicles used in the network. In particular, it restricts the number of pickup-space node arcs that depart from the origin depot with a remaining number of pickups equal to $L$. Constraints (30) and (31) define the domains of $X$ and $Y$.